Long-term debt ratio 0.2
Times interest earned 8.0
Current ratio 1.5
Quick ratio 1.0
Cash ratio 0.9
Inventory turnover 4.0
Average collection period 73 days

Use the above information from the tables to work out the following missing entries, and then calculate the company’s return on equity. Note: Turnover and the average collection period are calculated using start-of-year, not average, values. (Enter your answers in millions. Round intermediate calculations and final answers to 2 decimal places.)

INCOME STATEMENT
(Figures in $ millions)
  Net sales $   
  Cost of goods sold $   
  Selling, general, and administrative expenses 17.00   
  Depreciation 27.00   
Earnings before interest and taxes (EBIT) $  
  Interest expense $   
  Income before tax $   
  Tax (35% of income before tax) $   
  Net income $   
BALANCE SHEET
(Figures in $ millions)
This Year Last Year
  Assets
     Cash and marketable securities $    $ 27
     Accounts receivable $ 41
     Inventories $ 33
        Total current assets $    $ 101
     Net property, plant, and equipment $ 32
        Total assets $    $133
  Liabilities and shareholders’ equity
     Accounts payable $20.00 $ 15
     Notes payable 30.00 35
        Total current liabilities $ 50
     Long-term debt $ 27
     Shareholders’ equity $ 56
        Total liabilities and shareholders’ equity $195.00 $133
Answer #1

INCOME STATEMENT

(Figures in $ millions)

  Net sales

$205   

  Cost of goods sold

$132   

  Selling, general, and administrative expenses

17.00   

  Depreciation

27.00   

Earnings before interest and taxes (EBIT)

$29  

  Interest expense

$3.63   

  Income before tax

$25.37   

  Tax (35% of income before tax)

$8.88   

  Net income

$16.49

BALANCE SHEET

(Figures in $ millions)

This Year

Last Year

  Assets

     Cash and marketable securities

$45   

$ 27

     Accounts receivable

$5

41

     Inventories

$25

33

        Total current assets

$75   

$ 101

     Net property, plant, and equipment

$120

32

        Total assets

$195   

$133

  Liabilities and shareholders’ equity

     Accounts payable

$20.00

$ 15

     Notes payable

30.00

35

        Total current liabilities

$50.00

50

     Long-term debt

$29

27

     Shareholders’ equity

$116

56

        Total liabilities and shareholders’ equity

$195.00

$133

Explanations:-

Total current liabilities = $20.00 + 30.00 = $50.00

Total current assets = Current Liability * Current Ratio

                                  = $50 × 1.5 = $75.00

Cash = Current Liability * Cash Ratio

                        = $50 × 0.9 = $45.00

Accounts receivable + Cash = Current Liability * Quick Ratio

                                               = $50.00 × 1.0 = $50.00

Accounts receivable = $50.00 – Cash = $50.00 – 45.00 = $5.00

Inventories = $75.00 – 45.00 – 5.00 = $25.00

Total assets = Total liabilities and shareholders’ equity = $195.00

Net property, plant, equipment = $195.00 – 75.00 = $120.00

Sales = (365 / Average collection period) × Beginning receivables = (365 / 73) × $41 = $205.00

Cost of goods sold = Inventory turnover × Beginning inventory = 4.0 × $33 = $132.00

EBIT = $205.00 – 132.00 – 17.00 – 27.00 = $29

Interest = EBIT / Times interest earned = $29 / $8.0 = $3.63

Income before tax = EBIT - Interest expense = $29.00 – 3.63 = $25.37

Tax = Income before tax × .35 = $25.37 × .35 = $8.88

Net income = Income before tax – Tax = $25.37 – $8.88 = $16.49

Long-term debt + Equity = Total liabilities and equity – Total current liablities = $195.00 – $50 = $145.00

Long-term debt ratio = .2 = Long-term debt / (Long-term debt + Equity) = Long-term debt / $145.00; LTD = $29.00

Shareholders' equity = Total liabilities and equity – Total current liabilities - Long-term debt = $195.00 – 50.00 – 29.00 = $116.00